# Rethinking Cross-Attention for Infrared and Visible Image Fusion

### v1

@@ Abstract
The salient information of an infrared image and the abundant texture of a visible image can be fused to obtain a comprehensive image. As can be known, the current fusion methods based on Transformer techniques for infrared and visible (IV) images have exhibited promising performance. However, the attention mechanism of the previous Transformer-based methods was prone to extract common information from source images without considering the discrepancy information, which limited fusion performance. In this paper, by reevaluating the cross-attention mechanism, we propose an alternate Transformer fusion network (ATFuse) to fuse IV images. Our ATFuse consists of one discrepancy information injection module (DIIM) and two alternate common information injection modules (ACIIM). The DIIM is designed by modifying the vanilla cross-attention mechanism, which can promote the extraction of the discrepancy information of the source images. Meanwhile, the ACIIM is devised by alternately using the vanilla cross-attention mechanism, which can fully mine common information and integrate long dependencies. Moreover, the successful training of ATFuse is facilitated by a proposed segmented pixel loss function, which provides a good trade-off for texture detail and salient structure preservation. The qualitative and quantitative results on public datasets indicate our ATFFuse is effective and superior compared to other state-of-the-art methods.
Index Terms:
Image fusion, Transformer, cross-attention, discrepancy information, segmented pixel loss.
I Introduction
@@
As the single-modality image has limited interpretation ability, it is difficult to meet the subsequent requirements for understanding real scenarios. Therefore, developing an effective image fusion (IF) technology is urgently needed to assist people in gaining a deep understanding of images or tackling advanced computer vision tasks. Furthermore, image fusion (IF) can provide enhanced quality images for various applications, such as remote sensing monitoring [1], object detection[2, 3], medical analysis[4], and even military engineering[5]. Specifically, infrared sensors can capture thermal radiation information in poor weather conditions but have low imaging resolution and lack fine details. Contrarily, the visible image has abundant texture information, which can compensate for this deficiency. Therefore, the fusion of IV images can integrate respective merits to form an informative image comprising salient target information and abundant textual details. Recently, infrared and visible image fusion (IV-IF) has reached many satisfying achievements. The existing fusion methods can be made up of traditional methods[6] and deep learning-based ones[7].
@@
A generally accepted classification for traditional fusion methods is multi-scale transform (MST)-based methods[8], sparse representation (SR)-based methods[9], saliency-based methods[10, 11, 12], optimization-based methods[13, 14], and hybrid-based methods[15, 16]. Among the traditional fusion methods, feature extraction and feature fusion are the two crucial steps, most of which are based on handcrafted techniques. However, blindly selecting complex transformations or representations for feature extraction often leads to time-consuming and information loss. Moreover, some manually designed fusion strategies may reduce final fusion performance as these fusion strategies are not optimized for the corresponding generated features.
@@
Compared to traditional methods, recent advancements in deep learning have demonstrated superior computational efficiency and generalization. Existing proposed deep learning-based fusion methods are frequently implemented by leveraging either the convolutional neural network (CNN) techniques or Transformer techniques. Additionally, CNN-based methods can be further categorized into Autoencoder (AE)-based[17, 18, 19], end-to-end CNN-based[20, 21, 22, 23, 24], and GAN-based methods[25, 26, 27, 28].
@@
By contrast, the Transformer is equipped with self-attention and cross-attention mechanisms, which can effectively extract long-range context information and has exhibited excellent performance in many vision tasks apart from natural language processing [29, 30, 31, 32]. Recently, Transformer-based architecture has been introduced into the image fusion community [33, 34, 35, 36, 37]. These attempts have demonstrated long-range dependence, extracted by the self-attention mechanism and cross-attention mechanism, can further improve the quality of fusion outcomes [38, 39].
Despite the comparable results achieved by existing Transformer-based fusion, several challenges still exist in this domain. First, the attention mechanism is only used to capture common information of source images, while the discrepancy information has not been effectively separated and utilized. Second, a single Transformer module can not completely extract common information. The aforementioned issues significantly reduce fusion efficiency. Third, the existing pixel loss functions typically adopt a fixed approach, i.e., pixel maximum or weighted average, to guide the fusion process, which does not effectively preserve complete information.
To address the above-mentioned drawbacks, we devise an alternate Transformer fusion network (ATFuse) based on a modified cross-attention mechanism. For capturing the discrepancy information from two source images, we modify the cross-attention mechanism and propose a discrepancy information injection module (DIIM). To fully extract common information and integrate long dependencies of the source images, we design an alternate common information injection module (ACIIM). In addition, we propose a segmented pixel loss function that utilizes different constraints on pixel values to guide the fusion network. This proposed loss function can ensure the preservation of intensity information in fused images adequately.
@@
  • An end-to-end ATFuse network is proposed for the fusion of IV images. Extensive experiments on multiple datasets show that our proposed ATFuse method achieves good effect and generalization ability.
@@
  • A discrepancy information injection module (DIIM) is proposed based on the cross-attention mechanism. With this DIIM, the unique features of the source images can be explored respectively.
  • An alternate common information injection module (ACIIM) is applied to the proposed framework, in which common information is fully retained in the final result.
@@
  • A segmented pixel loss function composed of different pixel intensity constraints is designed to train ATFuse so that the preservation of texture details and brightness information can reach a good trade-off in the fused results.
@@
The structure of this paper is as follows. Section II surveys the related work. Section III states the proposed ATFuse framework. Section IV presents the experimental analysis and discussion. Section V summarizes this paper.
II RELATED WORK
In this section, we review the current CNN-based and Transformer-based IV-IF methods.
II-A CNN-based Image Fusion Methods
@@
Deep learning (DL) has been widely introduced into computer vision tasks such as image restoration[40], person re-identification[41],[42], and image super-resolution[43]. Attributed to the superior feature representation ability, DL-based IV-IF methods also have occupied the leading position in recent years. CNN was initially introduced into the field of IF by Liu et al.[44], using Siamese convolutional networks to generate weight maps. Li et al.[17] proposed a dense-connection AE-based approach, DenseFuse, to fuse IV images, avoiding information loss of deep features in the fused results. Jian et al. [19] designed an attention mechanism to fuse deep features based on a symmetric AE network, which enhanced the salient information of infrared images in the fused results. Li et al. [18] trained a learnable fusion network at two stages to further preserve image details and overcome the limitations of hand-crafted fusion strategies. Similarly, Ma et al.[45] proposed an end-to-end model that utilized a salient target mask to guide the network training to highlight the thermal information. To improve the model generalization, Xu et al.[24] developed a unified unsupervised fusion framework to deal with different fusion tasks.
@@
Most GAN-based fusion methods use an unsupervised manner to force the distribution of the fused result approach distributions of source images by a specific loss function. As a key milestone, Ma et al[25] first applied GAN to IV-IF. After that, various GAN-based variants have been proposed to boost the fusion performance. For example, Li et al. [46] utilized the coupled generative adversarial network for the image fusion task. Ma et al. [47] designed a DDcGAN, in which the high- and low-version of the fused image produced by a generator were used to deceive two discriminators respectively, to achieve the preservation of infrared information and detail information. To prevent the fused image from being biased to one of the source images, Ma et al. [48] adopted multiclassification constraints to achieve information balance.
II-B Transformer-based Fusion Methods
@@
Transformer was first proposed by Vaswani et al. [29] to address natural language processing issues and achieved remarkable success. Subsequently, Dosovitskiy et al. [49] designed visual Transformers to conduct an image classification task. Due to its self-attention mechanism that can capture long-range dependencies, the Transformer has been applied to many computer vision tasks, such as target detection[50, 51, 52], video inpainting[53, 54, 55], and image super-resolution[56, 57, 58]. In recent, some Transformer-based methods have been presented to handle IV-IF [59, 60, 61, 62].
@@
Vs et al. [34] utilized a spatial branch and a Transformer branch fusion strategies to merge local and global information, respectively. Wang et al. [63] proposed a residual Swin Transformer to extract the global features that further enhanced the representation ability of the previous Transformer in IV-IF tasks. In a similar manner, Ma et al. [39] integrated cross-domain learning with Swin Transformer to fully use the inter- and intra-domain contexts for fusion. To overcome hand-crafted fusion rules existing in Transformer-based fusion models, Tang et al. [37] proposed an end-to-end Transformer architecture to fuse IV images.
@@
Upon reviewing the previous Transformer-based fusion methods, we discovered certain limitations and rethought the usage of the cross-attention mechanism. Most studies based on Transformer structure merely grasp the global information of an image, aiming to compensate for the drawbacks of convolutional operations only capturing local information. To the best of our knowledge, both self-attention and cross-attention do not consider the discrepancy information between multimodalities in the application of IV-IF. Additionally, to fully exploit common information and long-range dependencies of source images, we adopt an alternating extraction approach depending on the original cross-attention mechanism.
III METHODOLOGY
@@
In this section, the overall framework of our proposed ATFuse is first introduced in Section III-A. Subsequently, in Section III-B, two modified cross-attention modules, DIIM and ACIIM, specifically designed for image fusion tasks are described. Finally, the loss function in our work is presented in Section III-C.
III-A Framework Overview
Figure 1: The overall framework of ATFuse.
@@ III-A Framework Overview
As illustrated in Fig. 1, the overall architecture of our ATFuse consists of a feature extraction module, a feature fusion module, and a feature reconstruction module. First, the source IV images are fed into the feature extraction module to extract shallow features and then transform them into patch embedding. Afterward, a feature fusion module is established for fusing discrepancy features and common features. Finally, a feature reconstruction is utilized to map the fused features into a composite image. Here we briefly introduce the proposed architecture.
@@
Feature extraction. Since the convolutional layers have stability and enhanced optimization ability [64], a convolutional layer with a kernel size of $3\times 3$ is still used to extract shallow local features in the feature extraction module. After that, there is a batch normalization (BN) layer and a HardSwish activation function. The feature extraction process can be formulated as:
@@
where $I_{ir}$ and $I_{vi}$ are source images. $F_{ir}^{sf}$ and $F_{vi}^{sf}$ denote the output features of shallow feature extraction $SF(\cdot)$. Subsequently, we use the patch embedding module $PE(\cdot)$ to transform the feature maps $F_{ir}^{sf}$ and $F_{vi}^{sf}$ to yield a series of token sequences $F_{ir}^{token}$ and $F_{vi}^{token}$, respectively. Then these tokens are fed to the subsequent DIIM and ACIIM. The process of the patch embedding can be expressed as:
${\left\{F_{ir}^{token},F_{vi}^{token}\right\}=\left\{PE(F_{ir}^{sf}),PE(F_{vi}^{sf})\right\}.}$ (2)
Feature fusion. The $F_{ir}^{token}$ and $F_{vi}^{token}$ are fed to feature fusion module $FF(\cdot)$ to generate the fused feature $F_{f}$. The feature fusion process can be formulated as:
${F_{f}=FF(F_{ir}^{token},F_{vi}^{token}),}$ (3)
where $FF(\cdot)$ contains a DIIM ($DIIM(\cdot)$) and a pair of ACIIMs ($ACIIM(\cdot)$), which are designed to further extract global dependencies features, i.e., discrepancy features and common features, respectively. It can be represented as:
$Z_{1}=DIIM(F_{ir}^{token},F_{vi}^{token}),$ (4)
$Z_{2}=ACIIM(F_{vi}^{token},Z_{1}),$
$Z_{3}=ACIIM(F_{ir}^{token},Z_{2}),$
$FF(F_{ir}^{token},F_{vi}^{token})=Z_{3}+Z_{1},$
where $Z_{1}$, $Z_{2}$, and $Z_{3}$ are the outputs of DIIM, the first ACIIM, and the second ACIIM respectively. A detailed description of DIIM and ACIIM is provided in Section 3.2.
Feature reconstruction. The fused feature $F_{f}$ is fed into an upsampling module $UP(\cdot)$ to resize back to the source images. Besides, we use the residual network-based refining module $RE(\cdot)$ to further recover the detail information of the fused image $I_{f}$. The feature reconstruction process can be expressed as:
@@
III-B DIIM and ACIIM.
Figure 2: The architecture of discrepancy information injection module (DIIM).

### v2

# Conceptualization of this study, Methodology, Software

@@ Abstract
The fusion of infrared and visible images is essential in remote sensing applications, as it combines the thermal information of infrared images with the detailed texture of visible images for more accurate analysis in tasks like environmental monitoring, target detection, and disaster management. The current fusion methods based on Transformer techniques for infrared and visible (IV) images have exhibited promising performance. However, the attention mechanism of the previous Transformer-based methods was prone to extract common information from source images without considering the discrepancy information, which limited fusion performance. In this paper, by reevaluating the cross-attention mechanism, we propose an alternate Transformer fusion network (ATFusion) to fuse IV images. Our ATFusion consists of one discrepancy information injection module (DIIM) and two alternate common information injection modules (ACIIM). The DIIM is designed by modifying the vanilla cross-attention mechanism, which can promote the extraction of the discrepancy information of the source images. Meanwhile, the ACIIM is devised by alternately using the vanilla cross-attention mechanism, which can fully mine common information and integrate long dependencies. Moreover, the successful training of ATFusion is facilitated by a proposed segmented pixel loss function, which provides a good trade-off for texture detail and salient structure preservation. The qualitative and quantitative results on public datasets indicate our ATFusion is effective and superior compared to other state-of-the-art methods.
keywords:
[…]
{highlights}
Proposal of ATFusion Network: An end-to-end ATFusion network is introduced for infrared and visible (IV) image fusion, demonstrating superior performance and strong generalization across multiple datasets.
Dual Cross-Attention Feature Injection: A novel discrepancy information injection module (DIIM) and alternate common information injection module (ACIIM) are proposed to effectively extract and integrate both distinctive and shared features from source images.
Segmented Pixel Loss for Balanced Fusion: A segmented pixel loss function with intensity-specific constraints is designed to optimize texture detail preservation and brightness balance in the fused results.
1 Introduction
@@
As the single-modality image has limited interpretation ability, it is difficult to meet the subsequent requirements for understanding real scenarios. Therefore, developing an effective image fusion (IF) technology is urgently needed to assist people in gaining a deep understanding of images or tackling advanced computer vision tasks. Furthermore, image fusion (IF) can provide enhanced quality images for various applications, such as infrared target detection[1], unmanned aerial vehicle (UAV) imaging[2], and environmental monitoring[3]. The resulting fused image merges thermal information with detailed texture, enhancing accuracy for remote sensing applications. Recent developments in VIF methods span traditional approaches[4] to advanced deep learning techniques[5, 6].
@@
A generally accepted classification for traditional fusion methods is as follows: multi-scale transform (MST)-based methods [7], sparse representation (SR)-based methods [8], saliency-based methods [9], optimization-based methods [10], and hybrid-based methods [11]. Among the traditional fusion methods, feature extraction and feature fusion are the two crucial steps, most of which are based on handcrafted techniques. However, blindly selecting complex transformations or representations for feature extraction often leads to time-consuming and information loss. Moreover, some manually designed fusion strategies may reduce final fusion performance as these fusion strategies are not optimized for the corresponding generated features.
@@
Compared to traditional methods, recent advancements in deep learning have demonstrated superior computational efficiency and generalization. Existing proposed deep learning-based fusion methods are frequently implemented by leveraging either the convolutional neural network (CNN) techniques or Transformer techniques. Additionally, CNN-based methods can be further categorized into Autoencoder (AE)-based [12], end-to-end CNN-based [13, 14], and GAN-based methods [15].
@@
By contrast, the Transformer is equipped with self-attention and cross-attention mechanisms, which can effectively extract long-range context information and has exhibited excellent performance in many vision tasks as the natural language processing [16]. Recently, Transformer-based architecture has been introduced into the image fusion community [17, 18]. These attempts have demonstrated long-range dependence, thanks to the self-attention and cross-attention mechanisms, further improving the quality of fusion outcomes [19].
Despite the comparable results achieved by existing Transformer-based fusion, several challenges still exist in this domain. First, the attention mechanism is only used to capture common information of source images, while the discrepancy information has not been effectively separated and utilized. Second, a single Transformer module cannot completely extract common information. The aforementioned issues significantly reduce fusion efficiency. Third, the existing pixel loss functions typically adopt a fixed approach, pixel maximum or weighted average, to guide the fusion process, which does not effectively preserve complete information.
To address the above-mentioned drawbacks, we devise an alternate Transformer fusion network (ATFusion) based on a modified cross-attention mechanism. For capturing the discrepancy information from two source images, we modify the cross-attention mechanism and propose a discrepancy information injection module (DIIM). To fully extract common information and integrate long dependencies of the source images, we design an alternate common information injection module (ACIIM). In addition, we propose a segmented pixel loss function that utilizes different constraints on pixel values to guide the fusion network. This proposed loss function can adequately ensure the preservation of intensity information in fused images.
@@
    An end-to-end ATFusion network is proposed for the fusion of IV images. Extensive experiments on multiple datasets show that our ATFusion method achieves good performance and generalization ability.
  2. 2.
@@
    A discrepancy information injection module (DIIM) and an alternate common information injection module (ACIIM) are proposed based on the cross-attention mechanism. With DIIM and ACIIM, the discrepancy and common features of the source images can be fully explored respectively.
  3. 3.
@@
    A segmented pixel loss function composed of different pixel intensity constraints is designed to train ATFusion so that the preservation of texture details and brightness information can reach a good trade-off in the fused results.
@@
The rest of this paper is organized as follows. Section II provides a survey of the related work. Section III presents the proposed ATFusion framework. Section IV details the experimental analysis with some discussions. Finally, Section V summarizes the key findings of this paper.
2 Related Work
In this section, we review the current CNN-based and Transformer-based VIF methods.
2.1 CNN-based Image Fusion Methods
@@
Deep learning (DL) has been widely introduced into computer vision tasks such as image restoration [20], person re-identification [21, 22], and image super-resolution [23]. Attributed to the superior feature representation ability, DL-based VIF methods have also occupied the leading position in recent years. CNN was initially introduced into the field of image fusion by Liu et al. [24], using Siamese convolutional networks to generate weight maps. Li et al. [25] proposed a dense-connection AE-based approach, DenseFuse, to fuse IV images, avoiding information loss of deep features in the fused results. Jian et al. [26] designed an attention mechanism to fuse deep features based on a symmetric AE network, which enhanced the salient information of infrared images in the fused results. Li et al. [27] trained a learnable fusion network at two stages to further preserve image details and overcome the limitations of hand-crafted fusion strategies. Similarly, Ma et al. [28] proposed an end-to-end model that utilized a salient target mask to guide the network training to highlight the thermal information. To improve the model generalization, Xu et al. [29] developed a unified unsupervised fusion framework to deal with different fusion tasks. To efficiently utilize the multi-scale deep features, Wang et al. [30] proposed a multi-scale densely connected encoder-decoder fusion architecture. Moreover, Wang et al. [31] integrated the dense Res2Net and attention into the AE model to consider the long-range dependency features in the fusion results.
@@
Most GAN-based fusion methods force, in an unsupervised manner, the distribution of the fused result by a specific loss function. As a key milestone, Ma et al. [32] first applied GAN to VIF. After that, various GAN-based variants have been proposed to boost the fusion performance. For example, Li et al. [33] utilized the coupled generative adversarial network for the image fusion task. Ma et al. [34] designed a DDcGAN, in which the high- and low-versions of the fused image produced by a generator were used to deceive two discriminators respectively, to achieve the preservation of infrared information and detail information. To prevent the fused image from being biased to one of the source images, Ma et al. [35] adopted multiclassification constraints to achieve information balance. Wang et al. [36] designed a cross-scale iterative attention manner to compute activity levels of different modal images based on the GAN model. To change the simple concatenation in generative adversarial fusion methods, Wang et al. [37] developed an interactive compensatory attention adversarial learning network.
2.2 Transformer-based Fusion Methods
@@
Transformer was first proposed by Vaswani et al. [38] to address natural language processing issues and achieved remarkable success. Subsequently, Dosovitskiy et al. [39] designed visual Transformers to conduct an image classification task. Due to its self-attention mechanism that can capture long-range dependencies, the Transformer has been applied to many computer vision tasks, such as target detection [40, 41], video inpainting [42, 43], and image super-resolution [44, 45]. Recently, some Transformer-based methods have been presented to handle VIF [46, 47].
@@
Vs et al. [48] utilized a spatial branch and a Transformer branch fusion strategies to merge local and global information, respectively. Wang et al. [49] proposed a residual Swin Transformer to extract the global features that further enhanced the representation ability of the previous Transformer in VIF tasks. In a similar manner, Ma et al. [19] integrated cross-domain learning with Swin Transformer to fully use the inter- and intra-domain contexts for fusion. To overcome hand-crafted fusion rules existing in Transformer-based fusion models, Tang et al. [18] proposed an end-to-end Transformer architecture to fuse IV images.
@@
Upon reviewing the previous Transformer-based fusion methods, we discovered certain limitations and rethought the usage of the cross-attention mechanism. Most studies based on Transformer structures merely grasp the global information of an image, aiming to compensate for the drawbacks of convolutional operations only capturing local information. To the best of our knowledge, both self-attention and cross-attention do not consider the discrepancy information between multimodalities when applied to VIF. Additionally, to fully exploit common information and long-range dependencies of source images, we adopt an alternating extraction approach depending on the original cross-attention mechanism.
3 Methodology
@@
In this section, we first introduce the overall framework of the proposed ATFusion, discussing its core design principles. Next, we detail the modified cross-attention modules, DIIM and ACIIM, which are specifically tailored for image fusion tasks. Finally, we describe the loss function used to optimize the network.
3.1 Framework Overview
Figure 1: The ATFusion framework: Feature Extraction for shallow feature and patch embedding, Feature Fusion via DIIM and ACIIM modules, and Feature Reconstruction with upsampling and refining to produce the fused image $I_{f}$.
@@ 3.1 Framework Overview
As illustrated in Fig. 1, the overall architecture of our ATFusion consists of a feature extraction module, a feature fusion module, and a feature reconstruction module. First, source IV images are fed into the feature extraction module to extract shallow features and then transform them into patch embedding. Afterward, a feature fusion module is established for fusing discrepancy features and common features. Finally, a feature reconstruction is utilized to map the fused features into a composite image. Here, we briefly introduce the proposed architecture.
@@
Feature extraction. Since convolutional layers have stability and improved optimization ability [50], a convolutional layer with a kernel size of $3\times 3$ is still used to extract shallow local features in the feature extraction module. After that, there is a batch normalization (BN) layer and a HardSwish activation function[51]. The feature extraction process can be formulated as:
@@
where $I_{ir}$ and $I_{vi}$ are the source images, and $F_{ir}^{sf}$ and $F_{vi}^{sf}$ denote the output features of the shallow feature extraction, $SF(\cdot)$. Subsequently, we use the patch embedding module, $PE(\cdot)$, to transform the feature maps $F_{ir}^{sf}$ and $F_{vi}^{sf}$ to yield a series of token sequences $F_{ir}^{t}$ and $F_{vi}^{t}$, respectively. Then, these tokens are fed to the subsequent DIIM and ACIIM. The process of the patch embedding can be expressed as:
${\left\{F_{ir}^{t},F_{vi}^{t}\right\}=\left\{PE(F_{ir}^{sf}),PE(F_{vi}^{sf})\right\}.}$ (2)
Feature fusion. The $F_{ir}^{t}$ and $F_{vi}^{t}$ are fed to the feature fusion module that is denoted as $FF(\cdot)$ to generate the fused feature $F_{f}$. The overall feature fusion process can be formulated as:
${F_{f}=FF(F_{ir}^{t},F_{vi}^{t})},$ (3)
where the feature fusion module contains a DIIM ($DIIM(\cdot)$) and a pair of ACIIMs ($ACIIM(\cdot)$), which are designed to further extract global dependencies features, discrepancy features and common features, respectively. Specifically, the DIIM utilizes the query vectors provided by $F_{vi}^{t}$ and the key-value vectors provided by $F_{ir}^{t}$ to fuse the discrepancy features between the source images. It can be represented as:
$Z_{1}=DIIM(F_{ir}^{t},F_{vi}^{t}),\$ (4)
where $Z_{1}$ represents the fused discrepancy features that contain the main structures of the source images. To inject detail information into the $Z_{1}$ features, the first ACIIM uses the key-value vectors provided by $F_{vi}^{t}$ to explore the common features mainly derived from the visible image.
$Z_{2}=ACIIM(F_{vi}^{t},Z_{1}),\$ (5)
where $Z_{2}$ indicates the fused features enhanced by the detail information of the visible image. Similarly, the $Z_{2}$ features still need to be further enhanced by alternately injecting detail information from the infrared image using the second ACIIM as:
$Z_{3}=ACIIM(F_{ir}^{t},Z_{2}),\$ (6)
where $Z_{3}$ are the fused features conveying the source images’ discrepancy information and common information. The alternative detail injection may weaken the information about the main structure of the fused features. Hence, we again add the $Z_{1}$ features for preservation:
${F_{f}}=Z_{3}+Z_{1}.$ (7)
A detailed description of DIIM and ACIIM is provided in Section 3.2.
Feature reconstruction. The fused feature $F_{f}$ is fed into an upsampling module, $UP(\cdot)$, to resize back to the source images. Besides, we use the residual network-based refining module, $RE(\cdot)$, to further recover the detail information of the fused image, $I_{f}$. The feature reconstruction process can be expressed as:
@@
3.2 DIIM and ACIIM